OpenStudy (anonymous):

Help Solve. 5x2 - 3x = (5x2 - 7x) + (8x - 2) A. x = -2 B. x = 2 C. x = -1/2 D. x = 1/2

OpenStudy (experimentx):

C

OpenStudy (anonymous):

how

OpenStudy (experimentx):

cancel you 5x^2, you have -3x = x - 2 => -4x = -2 => x = 1/2